Are these lines parallel
y=-7x-9
-3y=21x+7

Answer by engsaoudah(9)

y=-7x-9
-3y=21x+7
these two lines are parallel because they have the same gradient when you write the second equation in the form y=MX+C
-3y=21x+7
divide by -3
y=-7x-7/3
m=gradient x=x-axis C=y-intercept y=y-axis
m=-7 in both cases so the lines are parallel.